# The Atlas Model and SDEs with Boundary Interaction

### v1

@@ 1.3 Reformulation via McKean–Vlasov SDEs with Interaction Through Hitting and Local Times
To motivate the representation of PDE (1.8) through a reflected McKean–Vlasov SDE, let us first discuss its more canonical probabilistic representation in terms of a McKean–Vlasov SDE with interaction through hitting times [17, 26, 11]. This McKean–Vlasov SDE is given by
@@
for $x\in[0,\infty)$, where $\tau_{x}=\inf\{t>0\mathpunct{:}Y^{x}_{t}\leq 0\}$, and the initial condition $v$ is a signed measure. To connect this SDE with the mean-field limit (1.3), we must choose $v=\nu_{0}$. However, we intend to study McKean–Vlasov SDE (1.9) in more generality, where we do not assume that $v=\frac{1}{\alpha}-\mu_{0}$ for a probability measure $\mu_{0}$. Below we shall (implicitly) formulate an assumption on the signed measure $v$, generalising Assumption 1.6, that guarantees that the integral on the right-hand side of (1.10) is finite.
Formally, solutions to the above equation are defined in the following sense.
@@
Note that in the setting where $v$ is a probability measure, it is now well-established (cf. [11]) that McKean–Vlasov SDE (1.9) is indeed a probabilistic representation of the supercooled Stefan problem, PDE (1.8), in the following sense: if $\ell$ solves McKean–Vlasov SDE (1.9) and we define the distributions $\nu_{t}$, $t\geq 0$, on $\mathbb{R}$ by
@@
for Borel-measurable $A\subset[0,\infty)$ and $Y^{x}$ as in Equation (1.9), then $\nu_{t}$ is a solution to PDE (1.8) and $b_{t}=\alpha\ell_{t}$. Of course, some care must to be taken in formalising the precise sense in which $\nu_{t}$ solves PDE (1.8), since $\partial_{x}\nu_{t}(b_{t})$ may not exist for all times. For details, we refer to [11, Theorem 1.1]. In our analysis, we shall pass directly from McKean–Vlasov SDE (1.9) to the mean-field limit (1.3) of the Atlas model, circumventing the need for any regularity analysis of PDE (1.8). Let us note that if $\partial_{x}\nu_{t}(b_{t})$ does exist for all times and is integrable, then
@@
Note that when $m$ is derived from the density of the distribution of $\xi$, using Assumption 1.6, then it satisfies Assumption 1.10. However, if we only suppose that Assumption 1.10 holds, then the nonnegative function $x\mapsto(\frac{1}{\alpha}-m([0,x]))=(\frac{1}{\alpha}-v(x))$ need not even integrate to a finite value.
@@
$\int_{\mathbb{R}}\partial_{x}\nu_{t}(x)\varphi(x)\,\mathrm{d}x=\int_{[0,\infty)}\mathbb{E}[\varphi(X^{x}_{t}+\alpha\ell_{t})]\,\mathrm{d}m(x)$ (1.14)
for Borel measurable and bounded $\varphi\mathpunct{:}\mathbb{R}\to\mathbb{R}$. In the case where $m$ is a finite measure, this is explained in more detail in [3, Subsection 1.5.2]. Here, we shall exclusively focus on the probabilistic perspective and show that McKean–Vlasov SDEs (1.9) and (1.11) have the same set of solutions. This follows from the easily established equality
@@
The growth condition (1.13) is necessary to ensure that the integral on the right-hand side of (1.12) is finite (see Lemma 4.2). In that sense, the assumptions of Theorem 1.12 are essentially minimal for existence in the (sub)critical case. By critical (subcritical), we mean that the coefficient $\alpha$ appearing in (1.11) and the maximal mass $\sup_{x\geq 0}m([0,x])$ satisfy $\alpha\sup_{x\geq 0}m([0,x])\leq 1$ ($\alpha\sup_{x\geq 0}m([0,x])<1$). Theorem 1.12 shows that in the (sub)critical case, McKean–Vlasov SDE (1.11) has a continuous global solution. In the supercritical regime,
@@
solutions may only exist locally or exhibit jumps. A case where continuous but only local-in-time solutions exist is discussed in [3, Theorem 1.12] under the assumption that $m$ is a probability measure. The supercooled Stefan problem with nonnegative and integrable initial condition is a special case in which global solutions exist but jumps may occur [17, 26, 11].
The proof of Theorem 1.12 is based on an application of the Schauder fixed-point theorem. Since we are in the critical regime, the main challenge lies in finding a stable set $\mathcal{K}\subset C([0,\infty))$ for the fixed-point map whose image under the fixed-point map is contained in a compact subset of $\mathcal{K}$. Due to the criticality, one might worry that elements of $\mathcal{K}$ grow unboundedly under repeated application of the fixed-point map and, thus, escape from $\mathcal{K}$. To construct a suitable $\mathcal{K}$, we draw on the global solution to McKean–Vlasov SDE (1.11), obtained in Baker, Hambly & Jettkant [3] in the case that $m$ is a measure, as an upper bound.
@@ 1.4 Related Literature
There is a rich literature on ranked-based particle systems both with a finite and a countably infinite number of particles, dating back to works of Harris [18] and Sznitman [35] on rankings of Brownian motion. Much attention has been dedicated to the semi-infinite Atlas model, consisting of a countable collection of particles, started from the points of a Poisson process, that diffuse according to independent Brownian motions, with only the lowest ranked particle assigned a positive drift. Pal & Pitman [27] study the long-range behaviour of the spacings between particles, showing convergence to a stationary distribution given by independent exponential distributions. They also hypothesise that a suitable scaling of the $i$th ranked particle converges to a fractional Brownian motion with Hurst parameter $\frac{1}{4}$. The validity of this conjecture follows from results by Dembo & Tsai [13] who prove that rescaled fluctuations of the Atlas model around its equilibrium follow a stochastic heat equation with Neumann boundary condition.
@@
As we already alluded to above, Cabezas, Dembo, Sarantsev & Sidoravicius [7] analyse the large $N$ limit of the semi-infinite analogue of (1.1), where particles are started from a constant-rate Poisson process. Due to the simple initial profile, they can explicitly determine the system’s limiting density, which solves a Stefan-type moving boundary problem. Depending on whether the rate of the Poisson process lies within $(0,2\gamma)=(0,\frac{1}{\alpha})$ or is above $2\gamma=\frac{1}{\alpha}$, the deterministic limit of the Atlas particle, which plays the role of the moving boundary, follows a square-root trajectory with a positive or negative coefficient.
@@
The more recent work by Atar & Budhiraja [2], generalises [7] to varying initial configurations provided by locally finite measures. They characterise the limiting profile through a moving boundary problem involving measures that generalises the one-phase Stefan problem from [7]. In particular, the limit of the Atlas particle is no longer shown to follow a deterministic trajectory but is instead represented by a measure. This is analogous and served as an inspiration for our generalised formulation of McKean–Vlasov SDE (1.3) from Definition 1.3. Only if the initial configuration $\mu_{0}$ is sufficiently dense in the sense that $\mathrm{d}\mu_{0}(x)\geq\lambda\,\mathrm{d}x$ for some $\lambda>0$, a deterministic trajectory can be recovered. This result is obtained by exploiting the fact that the gaps of the particle system with the dense initial profile $\mu_{0}$ stochastically dominate those from the system with the constant initial density $\lambda$. Then one can proceed similarly to [7] to conclude. Note that our Assumption 1.6 does not allow for a comparison argument, so the proof techniques we employ are quite different from those in [2]. Let us emphasise though that our assumptions are not weaker than the ones from [2], so the results are (partially) complementary.
@@
Finite ranked-based particle systems were studied in much more generality than their infinite counterparts, with rank-dependence both in the drift and diffusion coefficients of all particles being considered. The first comprehensive existence result for weak solutions of such equations was provided by Bass & Pardoux [5], though their proper motivation comes from piecewise filtering. The applicability to particle systems with interaction through ranks is rather coincidental. Subsequently, Ichiba, Karatzas & Shkolnikov [20] showed that the weak solutions are in fact strong up until the first time three particles collide. Criteria for the absence and presence of such collision events were derived in [19, 20, 30]. Banner, Fernholz & Karatzas [4] study the long-term behaviour of generalised Atlas models with finitely many particles. Corresponding results for the simple semi-infinite Atlas model can be found in [27, 31].
@@
Several articles consider propagation of chaos for particle systems where the rank-dependence of the coefficients is expressed through a dependence on the empirical cumulative distribution function (CDF) of the particles. Note that this structural requirement precludes the Atlas model due to the exploding drift of lowest ranked particle. In many cases, the limit of the empirical CDF can be shown to satisfy a generalised one-dimensional porous medium equation. The solution of this equation describes the CDF of the representative particle in the mean-field limit. One of the earliest works in this direction is that of Jourdain [21], who studies particle approximations of the classical porous medium equation. Jourdain establishes propagation of chaos for the approximating ranked-based particle system and proves that the limit of the empirical CDF is indeed a solution of the classical porous medium equation. These results were extended to a broader range of coefficients by Jourdain & Reygner [22]. In a preceding work, Shkolnikov [33] proves propagation of chaos for initial configurations for which the gaps between consecutive particle ranks are stationary. Since then various extensions have appeared: large deviations [12], a central limit theorem [25], and the addition of common noise [24, 34].
@@
Lastly, let us comment on the connection between the Atlas model and Stefan-type moving boundary problems (see e.g. [8, 32, 14, 10] and references therein). As discussed earlier, this connection was already explored in [7] for initial densities with constant value $\lambda>0$. If $\lambda>2\gamma=\frac{1}{\alpha}$, the boundary in PDE (1.7) is nonincreasing, corresponding to the classical Stefan problem describing the melting of a frozen liquid. For $\lambda\in(0,2\gamma)$, the boundary advances into the domain, as is the case for the supercooled Stefan problem that models freezing of a supercooled liquid. For nonconstant initial conditions, treated in [2] and the present work, the boundary need not be monotonic, yielding a mixture of both models. While [7] and [2] work directly with PDE (1.7) (or in the latter case with an integrated and relaxed version introduced by Atar [1]), we proceed via the transformed PDE (1.8). The latter leads us to a precise relationship between the Atlas model and the supercooled Stefan problem from [17, 26], which admits the convenient probabilistic representation (1.9). Solutions to this representation can be constructed by extending ideas from Baker, Hambly & Jettkant [3].
@@ 1.5 Main Contributions and Structure of the Paper
Let us conclude this section with a discussion of our contributions and an outline of the paper. The first contribution we wish to highlight is our new probabilistic interpretation of the mean-field limit of the Atlas model as a novel type of reflected SDE with a constraint on the mean-growth of the regulator process. Additionally, we introduce a relaxed formulation of this SDE and rigorously connect it with various existing models.
@@
Section 3 serves a dual purpose. First, we show that any limit point of the particle system (1.1) is a solution to McKean–Vlasov SDE (1.3) in the generalised sense, thereby completing the proof of Theorem 1.4. Then, we prove the uniqueness result for generalised solutions, Theorem 1.5, drawing on approximations by solutions of (1.3) in the strong sense. Unlike [2], our arguments are probabilistic in nature, avoiding an analysis of PDEs (1.7) and (1.8).
@@ 2 The Finite Atlas Model
In this section, we analyse the particle system (1.1) through the lens of the reflected SDE (1.2) whose boundary is determined by a constraint on the growth of the empirical average of the regulator processes. We prove a comparison result for SDE (1.2), which, in particular, implies uniqueness. This is subsequently used to establish tightness of the particle system.
@@
Let $X^{1}$, …, $X^{N}$ be a solution to the Atlas model 1.1, guaranteed to exist by the theory of Bass & Pardoux [5]. Note that this solution is a priori weak, so we tacitly assume that $(\Omega,\mathcal{F},\mathbb{F}^{N},\mathbb{P})$ is sufficiently rich to support $X^{1}$, …, $X^{N}$. We already saw in the introduction that setting $B^{N}=X^{(1)}$ yields a solution to SDE (1.2). Moreover, the corresponding solutions of the Skorokhod problems for $\xi_{i}+W^{i}-B^{N}$ is given by $(X^{i}-B^{N},L^{i})$, where $L^{i}=(L^{i}_{t})_{t\geq 0}$ is defined as $L^{i}_{t}=\int_{0}^{t}\gamma N\mathbf{1}_{\{X^{i}_{s}=B^{N}_{s}\}}\,\mathrm{d}s$. According to the following comparison result, this is also the only solution to SDE (1.2).
@@
For $k=1$, 22, let $B^{k}$ be a solution to SDE (1.2) with $\mathcal{F}^{N}_{0}$-measurable initial conditions $\xi^{k}_{1}$, …, $\xi^{k}_{N}$ and denote the solution to the Skorokhod problem for $\xi^{k}_{i}+W^{i}-B^{k}$ by $(X^{k,i}-B^{k},L^{k,i})$, $i=1$, …, $N$. Then, if $\xi^{1}_{i}\leq\xi^{2}_{i}$ for $i\in[N]$, we have that $X^{1,i}_{t}\leq X^{2,i}_{t}$ and $B^{1}_{t}\leq B^{2}_{t}$ for $t\geq 0$ and $i\in[N]$. In particular, SDE (1.2) exhibits pathwise uniqueness.
@@
The proof is similar to that of Proposition 1.2, which concerns the mean-field limit. Let $B^{k}$, $X^{k,i}$, and $L^{k,i}$, $i\in[N]$ and $k\in\{1,2\}$ be as in the statement of the proposition. Then by the fundamental theorem of calculus, we have
@@
for $t\geq 0$. This concludes the proof. ∎

### v2

@@ 1.3 Reformulation via McKean–Vlasov SDEs with Interaction Through Hitting and Local Times
To motivate the representation of PDE (1.8) through a reflected McKean–Vlasov SDE, let us first discuss its more canonical probabilistic representation in terms of a McKean–Vlasov SDE with interaction through hitting times [undefp, undefy, undefj]. This McKean–Vlasov SDE is given by
@@
for $x\in[0,\infty)$, where $\tau_{x}=\inf\{t>0\mathpunct{:}Y^{x}_{t}\leq 0\}$, and the initial condition $v$ is a signed measure. To obtain a representation of PDE (1.8) in terms of this SDE, we must choose $v=\nu_{0}$. However, we intend to study McKean–Vlasov SDE (1.9) in more generality, so we do not necessarily assume that $v$ is of the form $\nu_{0}=\frac{1}{\alpha}-\mu_{0}$ for a probability measure $\mu_{0}$. Instead, we shall below (implicitly) formulate an assumption on the signed measure $v$, generalising Assumption 1.6, which guarantees that the integral on the right-hand side of (1.9) is finite.
Formally, solutions to Equation (1.9) are defined in the following sense.
@@
Note that in the setting where $v$ is a probability measure, it is now well-established (cf. [undefj]) that McKean–Vlasov SDE (1.9) is indeed a probabilistic representation of the supercooled Stefan problem, PDE (1.8), in the following sense: if $\ell$ solves McKean–Vlasov SDE (1.9) and we define the distributions $\nu_{t}$, $t\geq 0$, on $\mathbb{R}$ by
@@
for Borel-measurable $A\subset[0,\infty)$ and $Y^{x}$ as in Equation (1.9), then $\nu_{t}$ is a solution to PDE (1.8) and $b_{t}=\alpha\ell_{t}$. Of course, some care must be taken in formalising the precise sense in which $\nu_{t}$ solves PDE (1.8), since $\partial_{x}\nu_{t}(b_{t})$ may not exist for all times. For details, we refer to [undefj, Theorem 1.1]. In our analysis, we shall pass directly from McKean–Vlasov SDE (1.9) to the mean-field limit (1.3) of the Atlas model, circumventing the need for any regularity analysis of PDE (1.8). Let us note that if $\partial_{x}\nu_{t}(b_{t})$ does exist for all times and is integrable, then
@@
Here $\lvert m\rvert$ is the variation of $m$.
Note that when $m$ is derived from the density of the distribution of $\xi$ using Assumption 1.6, then it satisfies Assumption 1.11. However, if we only suppose that Assumption 1.11 holds, then the nonnegative function $x\mapsto(\frac{1}{\alpha}-m([0,x]))=(\frac{1}{\alpha}-v(x))$ need not even integrate to a finite value, so it does not necessarily determine a probability distribution.
@@
$\int_{A}\partial_{x}\nu_{t}(x)\,\mathrm{d}x=\int_{[0,\infty)}\mathbb{P}\bigl((X^{x}_{t}+\alpha\ell_{t})\in A\bigr)\,\mathrm{d}m(x)$ (1.14)
for Borel measurable $A\subset[0,\infty)$. In the case where $m$ is a finite measure, this is explained in more detail in [undefb, Subsection 1.5.2]. Here, we shall exclusively focus on the probabilistic perspective and show that McKean–Vlasov SDEs (1.9) and (1.11) have the same set of solutions. This follows from the easily established equality
@@
In the statement of the above proposition, it is understood that the initial condition $v$ of (1.9) and the initial condition $m$ of (1.11) are related by $v(x)=m([0,x])$ for $x\geq 0$.
@@
The growth condition (1.13) is necessary to ensure that the integral on the right-hand side of (1.12) is finite (see Lemma 4.2). In that sense, the assumptions of Theorem 1.13 are essentially minimal for existence in the (sub)critical case. Recall that by critical (subcritical) we mean that the coefficient $\alpha$ appearing in (1.11) and the maximal mass $\sup_{x\geq 0}m([0,x])$ satisfy $\alpha\sup_{x\geq 0}m([0,x])\leq 1$ ($\alpha\sup_{x\geq 0}m([0,x])<1$). Theorem 1.13 shows that in the (sub)critical case, McKean–Vlasov SDE (1.11) has a continuous global solution. In the supercritical regime,
@@
solutions may only exist locally or exhibit jumps. A case where continuous but only local-in-time solutions exist is discussed in [undefb, Theorem 1.12] under the assumption that $m$ is a probability measure. The supercooled Stefan problem with nonnegative and integrable initial condition is a special case in which global solutions exist but jumps may occur [undefp, undefy, undefj].
The proof of Theorem 1.13 is based on an application of the Schauder fixed-point theorem. Since we are in the critical regime, the main challenge lies in finding a stable set $\mathcal{K}\subset C([0,\infty))$ for the fixed-point map whose image under the fixed-point map is contained in a compact subset of $\mathcal{K}$. Due to the criticality, one might worry that elements of $\mathcal{K}$ grow unboundedly under repeated application of the fixed-point map and, thus, escape from $\mathcal{K}$. To construct a suitable $\mathcal{K}$, we employ as an upper bound the global solution to McKean–Vlasov SDE (1.11) obtained in Baker, Hambly & Jettkant [undefb] for initial conditions given by finite measures. The applicability of this theory crucially rests on the fact that $m(\{0\})<\frac{1}{\alpha}$ and $m([0,x])\leq\frac{1}{\alpha}$ for $x\geq 0$.
@@ 1.4 Related Literature
There is a rich literature on ranked-based particle systems both with a finite and a countably infinite number of particles, dating back to works of Harris [undefq] and Sznitman [undefah] on rankings of Brownian motion. Much attention has been dedicated to the semi-infinite Atlas model, consisting of a countable collection of particles, started from the points of a Poisson process, that diffuse according to independent Brownian motions, with only the lowest ranked particle assigned a positive drift. Pal & Pitman [undefz] study the long-range behaviour of the spacings between particles, showing convergence to a stationary distribution given by independent exponential distributions. They also hypothesise that a suitable scaling of the $i$th ranked particle converges to a fractional Brownian motion with Hurst parameter $\frac{1}{4}$. The validity of this conjecture follows from results by Dembo & Tsai [undefl] who prove that rescaled fluctuations of the Atlas model around its equilibrium follow a stochastic heat equation with Neumann boundary condition.
@@
As we already alluded to above, Cabezas, Dembo, Sarantsev & Sidoravicius [undeff] analyse the large $N$ limit of the semi-infinite analogue of (1.1), where particles are started from a constant-rate Poisson process. Due to the simple initial profile, they can explicitly determine the system’s limiting density, which solves a Stefan-type moving boundary problem. Depending on whether the rate of the Poisson process lies within $(0,2\gamma)=(0,\frac{1}{\alpha})$ or is above $2\gamma=\frac{1}{\alpha}$, the deterministic limit of the Atlas particle, which plays the role of the moving boundary, follows a square-root trajectory with a positive or negative coefficient.
@@
The more recent work by Atar & Budhiraja [undefa], generalises [undeff] to varying initial configurations provided by locally finite measures. They characterise the limiting profile through a moving boundary problem involving measures that generalises the one-phase Stefan problem from [undeff]. In particular, the limit of the Atlas particle is no longer shown to follow a deterministic trajectory but is instead represented by a measure. This is analogous and served as an inspiration for our generalised formulation of McKean–Vlasov SDE (1.3) from Definition 1.3. Only if the initial configuration $\mu_{0}$ is sufficiently dense in the sense that $\mathrm{d}\mu_{0}(x)\geq\lambda\,\mathrm{d}x$ for some $\lambda>0$, a deterministic trajectory can be recovered. This result is obtained by exploiting the fact that the gaps of the particle system with the dense initial profile $\mu_{0}$ are stochastically dominated by those from the system with the constant initial density $\lambda$. Then, one can proceed similarly to [undeff] to conclude. Note that our Assumption 1.6 does not allow for a comparison argument, so the proof techniques we employ are quite different from those in [undefa]. Let us emphasise though that our assumptions are not weaker than the ones from [undefa], so the results are (partially) complementary.
@@
Finite ranked-based particle systems were studied in much more generality than their infinite counterparts, with rank-dependence both in the drift and diffusion coefficients of all particles being considered. The first comprehensive existence result for weak solutions of such equations was provided by Bass & Pardoux [undefd], though their proper motivation comes from piecewise filtering. The applicability to particle systems with interaction through ranks is rather coincidental. Subsequently, Ichiba, Karatzas & Shkolnikov [undefs] showed that the weak solutions are in fact strong up until the first time three particles collide. Criteria for the absence and presence of such collision events were derived in [undefr, undefs, undefac]. Banner, Fernholz & Karatzas [undefc] study the long-term behaviour of generalised Atlas models with finitely many particles. Corresponding results for the simple semi-infinite Atlas model can be found in [undefz, undefad].
@@
Several articles consider propagation of chaos for particle systems where the rank-dependence of the coefficients is expressed through a dependence on the empirical cumulative distribution function (CDF) of the particles. Note that this structural requirement precludes the Atlas model due to the exploding drift of lowest ranked particle. In many cases, the limit of the empirical CDF can be shown to satisfy a generalised one-dimensional porous medium equation. The solution of this equation describes the CDF of the representative particle in the mean-field limit. One of the earliest works in this direction is that of Jourdain [undeft], who studies particle approximations of the classical porous medium equation. Jourdain establishes propagation of chaos for the approximating ranked-based particle system and proves that the limit of the empirical CDF is indeed a solution of the classical porous medium equation. These results were extended to a broader range of coefficients by Jourdain & Reygner [undefu]. In a preceding work, Shkolnikov [undefaf] proves propagation of chaos for initial configurations for which the gaps between consecutive particle ranks are stationary. Since then various extensions have appeared: large deviations [undefk], a central limit theorem [undefx], and the addition of common noise [undefw, undefag].
@@
Lastly, let us comment on the connection between the Atlas model and Stefan-type moving boundary problems (see e.g. [undefg, undefae, undefm, undefi] and references therein). As discussed earlier, this connection was already explored in [undeff] for initial densities with constant value $\lambda>0$. If $\lambda>2\gamma=\frac{1}{\alpha}$, the boundary in PDE (1.7) is nonincreasing, corresponding to the classical Stefan problem describing the melting of a frozen liquid. For $\lambda\in(0,2\gamma)$, the boundary advances into the domain, as is the case for the supercooled Stefan problem that models freezing of a supercooled liquid. For nonconstant initial conditions, treated in [undefa] and the present work, the boundary need not be monotonic, yielding a mixture of both models. While [undeff] and [undefa] work directly with PDE (1.7) (or in the latter case with an integrated and relaxed version introduced by Atar [undef]), we proceed via the transformed PDE (1.8). The latter leads us to a precise relationship between the Atlas model and the supercooled Stefan problem from [undefp, undefy], which admits the convenient probabilistic representation (1.9). Solutions to this representation can be constructed by extending ideas from Baker, Hambly & Jettkant [undefb].
@@ 1.5 Main Contributions and Structure of the Paper
Let us conclude this section with a discussion of our contributions and an outline of the paper. The first contribution we wish to highlight is our new probabilistic interpretation of the mean-field limit of the Atlas model as a novel type of reflected SDE with a constraint on the mean-growth of the regulator process. We introduce a relaxed formulation of this SDE and rigorously connect it with various existing models.
@@
Section 3 serves a dual purpose. First, we show that any limit point of the particle system (1.1) is a solution to McKean–Vlasov SDE (1.3) in the generalised sense, thereby completing the proof of Theorem 1.4. Then, we prove the uniqueness result for generalised solutions, Theorem 1.5, drawing on approximations by solutions of (1.3) in the strong sense. Unlike [undefa], our arguments are probabilistic in nature, avoiding an analysis of PDEs (1.7) and (1.8).
@@ 2 The Finite Atlas Model
In this section, we analyse the particle system (1.1) through the lens of the reflected SDE (1.2) whose boundary is determined by a constraint on the growth of the empirical average of the regulator processes. We prove several comparison results for SDE (1.2), which, in particular, imply uniqueness. These are subsequently used to establish tightness of the particle system.
@@
Let $X^{1}$, …, $X^{N}$ be a solution to the Atlas model (1.1), guaranteed to exist by the theory of Bass & Pardoux [undefd]. Note that this solution is a priori weak, so we tacitly assume that $(\Omega,\mathcal{F},\mathbb{F}^{N},\mathbb{P})$ is sufficiently rich to support $X^{1}$, …, $X^{N}$. We already saw in the introduction that setting $B^{N}=X^{(1)}$ yields a solution to SDE (1.2). Moreover, the corresponding solutions of the Skorokhod problems for $\xi_{i}+W^{i}-B^{N}$ is given by $(X^{i}-B^{N},L^{i})$, where $L^{i}=(L^{i}_{t})_{t\geq 0}$ is defined as $L^{i}_{t}=\int_{0}^{t}\gamma N\mathbf{1}_{\{X^{i}_{s}=B^{N}_{s}\}}\,\mathrm{d}s$. According to the following comparison result, this is also the only solution to SDE (1.2).
@@
For $k=1$, 22, let $N_{k}\geq 1$ and let $B^{k}$ be a solution to SDE (1.2) with $\mathcal{F}^{N_{k}}_{0}$-measurable initial conditions $\xi^{k}_{1}$, …, $\xi^{k}_{N_{k}}$ and denote the solution to the Skorokhod problem for $\xi^{k}_{i}+W^{i}-B^{k}$ by $(X^{k,i}-B^{k},L^{k,i})$, $i=1$, …, $N_{k}$.
  1. (i)
    If $N_{1}=N_{2}$ and $\xi^{1}_{i}\leq\xi^{2}_{i}$ for $i\in[N_{1}]$, then we have that $X^{1,i}_{t}\leq X^{2,i}_{t}$ and $B^{1}_{t}\leq B^{2}_{t}$ for $t\geq 0$ and $i\in[N_{1}]$. In particular, SDE (1.2) exhibits pathwise uniqueness.
  2. (ii)
    If $N_{1}=N_{2}$, $\xi^{1}_{i}\leq\xi^{2}_{i}$ for some $i\in[N_{1}]$, and $\xi^{1}_{j}=\xi^{2}_{j}$ for $j\in[N_{1}]\setminus\{i\}$, then $L^{1,i}_{t}\geq L^{2,i}_{t}$ for $t\geq 0$.
  3. (iii)
    If $N_{1}\leq N_{2}$ and $\xi^{1}_{i}=\xi^{2}_{i}$ for $i\in[N_{1}]$, then we have that $L^{1,i}_{t}\geq L^{2,i}_{t}$ for $t\geq 0$ and $i\in[N_{1}]$.
@@
We begin by establishing (i). The proof is similar to that of Proposition 1.2, which concerns the mean-field limit. For notational simplicity, set $N=N_{1}=N_{2}$ and let $B^{k}$, $X^{k,i}$, and $L^{k,i}$, $i\in[N]$ and $k\in\{1,2\}$ be as in the statement of the proposition. Then by the fundamental theorem of calculus, we have
@@
for $t\geq 0$. This concludes the proof of (i).
Let us show (ii) next, again setting $N=N_{1}=N_{2}$. Similarly to (2), we have for $k\geq 1$ and $t\geq 0$ that
$(X^{2,j}_{t}-X^{1,j}_{t})^{2k}$ $=(\xi^{2}_{j}-\xi^{1}_{j})^{2k}+2k\int_{0}^{t}(X^{2,j}_{s}-X^{1,j}_{s})^{2k-1}\,\mathrm{d}(L^{2,j}_{s}-L^{1,j}_{s})$
$\leq(\xi^{2}_{j}-\xi^{1}_{j})^{2k}+2k\int_{0}^{t}(B^{2}_{s}-B^{1}_{s})^{2k-1}\,\mathrm{d}(L^{2,j}_{s}-L^{1,j}_{s})$
We sum this inequality over $j\in[N]$ and then raise both sides to the power $\frac{1}{2k}$ to obtain
$\biggl(\sum_{j=1}^{N}(X^{2,j}_{t}-X^{1,j}_{t})^{2k}\biggr)^{\frac{1}{2k}}\leq\biggl(\sum_{j=1}^{N}(\xi^{2}_{j}-\xi^{1}_{j})^{2k}\biggr)^{\frac{1}{2k}}$
Letting $k\to\infty$ in both expressions yields
$X^{2,i}_{t}-X^{1,i}_{t}\leq\max_{j\in[N]}\lvert X^{2,j}_{t}-X^{1,j}_{t}\rvert\leq\max_{j\in[N]}|\xi^{2}_{j}-\xi^{1}_{j}\rvert=\xi^{2}_{i}-\xi^{1}_{i}.$
From this we deduce that
$\xi^{2}_{i}-\xi^{1}_{i}\geq X^{2,i}_{t}-X^{1,i}_{t}=\xi^{2}_{i}-\xi^{1}_{i}+L^{2,i}_{t}-L^{1,i}_{t},$
which finally implies that $L^{1,i}_{t}\geq L^{2,i}_{t}$.
It remains to prove the last statement of the proposition. For $K\geq 1$, we introduce an auxiliary system with initial conditions $\tilde{\xi}^{K}_{i}$, $i\in[N_{2}]$, defined by $\tilde{\xi}^{K}_{i}=\xi^{2}_{i}$ for $i\in[N_{1}]$ and $\tilde{\xi}^{K}_{i}=\xi^{2}_{i}\lor K$ for $i\in[N_{2}]\setminus[N_{1}]$. We denote the corresponding solution to SDE (1.2) by $\tilde{B}^{K}$ and let $(\tilde{X}^{K,i}-\tilde{B}^{K},\tilde{L}^{K,i})$ solve the Skorokhod problem for $\tilde{\xi}^{K}_{i}+W^{i}-\tilde{B}^{K}$. Since $\tilde{\xi}^{K}_{i}\geq\xi^{2}_{i}$ for $i\in[N_{2}]$, it holds by (i) that $\tilde{X}^{K,i}_{t}\geq X^{2,i}_{t}$ for $t\geq 0$. But the initial conditions of the particles in $N_{1}]$ coincide for the two systems of size $N_{2}$, so from $\tilde{X}^{K,i}_{t}\geq X^{2,i}_{t}$ we deduce that $\tilde{L}^{K,i}_{t}\geq L^{2,i}_{t}$ for $i\in[N_{1}]$.
Let us now define $\varrho_{K}$ to be the first time $t\geq 0$ such that $\tilde{X}^{K,i}_{t}=\tilde{B}^{K}_{t}$ for some $i\in[N_{2}]\setminus[N_{1}]$. By pathwise uniqueness of SDE (1.2), we have that $(\tilde{X}^{K,i}_{t},\tilde{L}^{K,i}_{t})=(X^{1,i}_{t},L^{1,i}_{t})$ for $t\in[0,\varrho_{K})$. Thus, if we can show that $\varrho_{K}\to\infty$ a.s. as $K\to\infty$, then it follows that
$L^{1,i}_{t}=\lim_{K\to\infty}L^{1,i}_{t\land\varrho_{K}}=\lim_{K\to\infty}\tilde{L}^{K,i}_{t\land\varrho_{K}}\geq\lim_{K\to\infty}L^{2,i}_{t\land\varrho_{K}}=L^{2,i}_{t}$
for $t\geq 0$ and $i\in[N_{1}]$. To see that $\varrho_{K}\to\infty$, we again appeal to the uniqueness of SDE (1.2), whereby $\tilde{B}^{K}_{t}=B^{1}_{t}$ for $t\in[0,\varrho_{K})$. Thus, on
$\{\varrho_{K}>0\}=\Bigl\{K>\min_{j\in[N_{1}]}\xi^{1}_{j}\Bigr\},$
we can alternatively write $\varrho_{K}$ as the first time $t\geq 0$ that $\tilde{X}^{K,i}=B^{1}_{t}$ for some $i\in[N_{2}]\setminus[N_{1}]$. Now, the process $B^{1}$ is a.s. locally bounded and, for $i\in[N_{2}]\setminus[N_{1}]$, we have $\tilde{X}^{K,i}_{t}\geq K+W^{i}_{t}$ for $t\geq 0$, so the first time that $\tilde{X}^{K,i}$ and $B^{1}$ meet indeed diverges as $K\to\infty$. This concludes the proof. ∎